The free fall acceleration on Mars is about one-third that on Earth. What If you throw a ball upward with the same velocity on Mars as on Earth:
a. How would you compare the balls maximum height to that on Earth?
b. How would its flight time compare?
1
Expert's answer
2021-10-06T15:52:40-0400

a. Find the maximum height on Earth and Mars:


"h_E=\frac{v^2}{2g},\space h_M=\frac{v^2}{2g/3},\\\space\\
\frac{h_M}{h_E}=3."

Therefore, the height on Mars is three times greater.

b. Flight time:


"t_E=\frac vg,\space t_M=\frac v{g/3},\\\space\\
\frac{t_M}{t_E}=3."

The same about time.